OpenStudy (alesia/bailey):

Please help if you know inequalities. -4<3w+5

OpenStudy (texaschic101):

-4 < 3w + 5 <= 20 ---- subtract 5 -4 - 5 < 3w + 5 - 5 <= 20 - 5 -- combine terms -9 < 3w <= 15 -- now divide by 3 -9/3 < 3w/3 <= 15/3 -3 < w <= 5 any questions ?

OpenStudy (alesia/bailey):

brb. im at school and its time to go.

OpenStudy (rabya!):

-4<3w+5 and 3w+5<or = 20 u can do by this method too !!